A solution of sodium hydroxide (NaOH) is prepared by taking 20.0 g NaOH and dissolving in enough water to make 500.0 mL of solution. What is its concentration in “molar” units? (Formula weight of NaOH is 40.0 g/mol.)

How many moles of NaOH are contained in 12.71 mL of this sodium hydroxide solution?
Answer #1

molarity = (W/MW) (1000 / V inmL)

W = 20.0g

MW = 40.0 g/mol

V = 500 mL

Molarity = (20.0 / 40) (1000 / 500)

Molarity = 1.0 M

now

molarity = n /V in L

n = Molarity x V inL

n = 1.0 x 12.71 / 1000

n = 0.0127 moles

moles of NaOH contain in 12.71 mL solution = 0.0127